### 28 Apr 2026 v1

# OxyGent: Making Multi-Agent Systems Modular, Observable, and Evolvable via Oxy Abstraction

Abstract
Deploying production-ready multi-agent systems (MAS) in complex industrial environments remains challenging due to limitations in scalability, observability, and autonomous evolution. We present OxyGent, an open-source framework that enables modular, observable, and evolvable MAS via a unified Oxy abstraction, in which agents, tools, LLMs, and reasoning flows are encapsulated as pluggable atomic components. This Lego-like assembly paradigm supports scalable system composition and non-intrusive monitoring. To enhance observability, OxyGent introduces permission-driven dynamic planning that replaces rigid workflows with execution graphs generated at runtime, which provide adaptive visualizations. To support continuous evolution, the framework integrates OxyBank, an AI asset management platform that supports automated data backflow, annotation, and joint evolution. Empirical evaluations and real-world case studies show that OxyGent provides a robust and scalable foundation for MAS. OxyGent is publicly available at https://oxygent.jd.com/.
OxyGent: Making Multi-Agent Systems Modular, Observable, and Evolvable via Oxy Abstraction
Junxing Huthanks: Equal contribution.   Tianlong Li11footnotemark: 1   Lei Yu   Ai Hanthanks: Corresponding author. JD.com Beijing, China junxing.hu@cripac.ia.ac.cn, {litianlong26, yulei116, hanai5}@jd.com
1 Introduction
Large Language Model (LLM) agents are transitioning from experimental prototypes to collaborative Multi-Agent Systems (MAS) capable of solving sophisticated real-world tasks Wang et al. (2024). However, in large-scale distributed environments such as intelligent assistants Fu et al. (2024), existing MAS frameworks encounter critical bottlenecks. Traditional development paradigms may lack standardized abstractions, leading to low efficiency and repetitive implementation of core functions. Moreover, rigid plan-and-execute workflows are too brittle to handle the uncertainties of dynamic environments. In addition, without a closed-loop mechanism to evaluate and refine agents using online feedback, performance tends to stagnate or even deteriorate Gao et al. (2025).
Figure 1: A multi-agent file management assistant built with OxyGent. (a) MAS visualization. (b) Permission relationships are defined in the implementation code.
To enhance scalability and development efficiency, OxyGent introduces a unified atomic abstraction in which every functional element, including agents, tools, and reasoning flows, is encapsulated as a pluggable Oxy node. Figure 1 shows an example of a file management assistant quickly built based on it. This Lego-like approach empowers researchers to rapidly assemble and hot-swap specialized components across diverse scenarios without complex manual reconfiguration. By integrating Aspect-Oriented Programming (AOP) Kiczales et al. (1997), the framework separates core business logic from cross-cutting concerns like logging and security auditing. This design ensures that as the number of agents grows, the system maintains a clean and modular architecture, facilitating seamless horizontal expansion in production settings.
To address the challenge of observability, OxyGent introduces a permission-driven dynamic planning mechanism. Unlike conventional frameworks that rely on static, pre-defined Directed Acyclic Graphs (DAGs), OxyGent defines the potential collaboration space through permission relationships between Oxy nodes, as shown in Figure 1(b). During inference, the system automatically synthesizes the execution path and generates real-time visualizations of the actual call graphs (Figure 1(a)). This transparency allows users to inspect and trace the decision-making trajectories of MAS, ensuring that even in volatile task environments, emergent behaviors remain monitorable and manageable.
To facilitate the agent evolution, OxyGent introduces OxyBank, a specialized platform for AI asset management. OxyBank serves as the evolutionary engine of OxyGent by providing a systematic pipeline that captures online execution traces and converts them into high-quality training samples. By integrating automated rewarding and human-in-the-loop annotation, OxyBank enables the backflow of domain-specific knowledge into the multi-agent system. This closed-loop process supports the joint evolution of agent strategies, ensuring that the collective intelligence of the ecosystem continuously improves through data-centric learning rather than remaining static after deployment. The main contributions are summarized as follows:
  • We propose the unified Oxy abstraction and an AOP-infused orchestration paradigm, enabling Lego-like construction and high scalability for complex multi-agent systems.
  • We implement permission-driven dynamic planning with real-time execution visualization, enabling deep observability into agent behaviors.
  • We introduce OxyBank, a platform for AI asset management that supports agent evolution through automated data backflow and evaluation.
  • Extensive evaluations and case studies demonstrate the effectiveness of OxyGent, proving its robustness in solving collaborative tasks.
2 Related Work
Orchestration Paradigms and Dynamic Planning. Current MAS frameworks focus on balancing agent autonomy with structural control through diverse paradigms. LangGraph LangChain AI (2024) serves as a low-level orchestration framework and runtime for building long-running, stateful agents via graph structures. CrewAI CrewAIInc (2025) utilizes a role-driven approach, combining the collaborative intelligence of “Crews” with the precise control of “Flows” to manage complex processes. Beyond graph and role-based designs, Semantic Kernel Microsoft (2024) integrates AI capabilities with enterprise business logic, while MetaGPT Hong et al. (2024) encodes Standardized Operating Procedures (SOPs) into prompt sequences to mimic software company workflows. Recent research has also explored more fluid architectures. Aime Shi et al. (2025) replaces fixed planners with dynamic actor instantiation, and AFlow Zhang et al. (2025) automates the generation of agentic workflows. OxyGent distinguishes itself by unifying agents, tools, and flows into interchangeable atomic nodes governed by permission-driven dynamic planning and real-time execution visualization.
Execution Lifecycle and System Observability. Observability is a critical production-readiness requirement, yet many frameworks prioritize developer experience over runtime monitoring. OpenAI Agents SDK OpenAI (2025), the successor to Swarm OpenAI (2024), adopts a minimalist set of primitives, including Agents, Handoffs, and Guardrails, and provides built-in tracing for debugging agentic flows. AutoGen Wu et al. (2024) models interactions as asynchronous message passing between entities, while Pydantic AI Pydantic (2024) introduces a type-safe Python framework that enforces strict contracts for tool calling and data exchange. For enterprise-grade monitoring, Strands Agents AWS (2025) provides a model-agnostic toolkit optimized for OpenTelemetry integration, similar to the durable execution principles in Dapr Agents Dapr (2025). OxyGent implements a comprehensive AOP-based lifecycle management system that modularly injects cross-cutting concerns such as security auditing and performance monitoring into distributed agent joinpoints.
Figure 2: The OxyGent framework provides four data scopes: Application, Session Group, Request, and Node. This multi-level data isolation and sharing ensures data management efficiency and development convenience in MAS.
Figure 3: The execution lifecycle of Oxy. A series of management and coordination steps ensures data flow and processing, enabling MAS to dynamically plan and generate flowcharts in real time.
Evolutionary AI Asset Management. The transition toward agents that learn and improve has led to systems that prioritize memory and training loops Han et al. (2025). Agno Agno-agi (2025) serves as specialized infrastructure for reasoning systems, utilizing its AgentOS layer to manage long-term memory and knowledge accumulation across sessions. AWorld Yu et al. (2025) provides a high-performance interaction runtime designed for large-scale experience generation to enable practical reinforcement learning for agents. EvoAgent Yuan et al. (2025) and SE-Agent Lin et al. (2025) have proposed automatic multi-agent extension via evolutionary operators and cross-trajectory refinement, respectively. Moreover, OWL Hu et al. (2025) improves cross-domain generalization by optimizing a domain-agnostic planner with reinforcement learning from real-world feedback, whereas Chain-of-Agents Zhang et al. (2024) focuses on distilling multi-agent reasoning patterns into unified foundation models. OxyGent facilitates a continuous build-inference-evolution cycle through OxyBank, an industrial-scale AI asset platform that unifies data annotation and knowledge backflow to improve collective agents.
Figure 4: The file management assistant is built on OxyGent. From left to right: MAS visualization, question-answering box, and the output of each node. As shown in Figure 10, each intermediate step of the reasoning process can be rolled back and reviewed.
Figure 5: MAS inference monitoring. OxyGent has built-in production-grade time tracking, which displays task distribution and resource congestion in real time, facilitating MAS architecture optimization. More features are introduced in Appendix A.1.
Figure 6: OxyBank, as OxyGent’s one-stop AI asset management platform, supports knowledge base construction, data annotation, memory management, multi-agent system evaluation, and enables agent evolution.
Figure 7: On July 22, 2025, OxyGent achieved the second-highest score (59.14%) among open-source methods on the GAIA leaderboard, second only to OWL++, which was the strongest open-source method at the time.
3 OxyGent
OxyGent aims to transform multi-agent orchestration from static, hard-coded workflows into an elastic, Lego-like assembly of atomic components. This section details the framework’s infrastructure, the dynamic planning mechanisms, and the evolutionary loop powered by OxyBank.
3.1 Unified Oxy Abstraction and Hierarchical Data Scopes
To facilitate the flexible construction of complex MAS, OxyGent introduces a unified atomic abstraction where agents, tools, LLMs, and flows are wrapped into pluggable Oxy components. Unlike traditional frameworks that treat these entities as distinct software layers, OxyGent provides a consistent interface for all AI atoms, enabling hot-swapping of reasoning logic and capabilities at runtime without refactoring.
A cornerstone of unified abstraction is the four-tier data scoping mechanism, which provides a structured approach to state management in distributed environments (Figure 2). Rather than relying on a single global state, OxyGent isolates data into specific domains to prevent data-privilege leaks and ensure efficiency. The domain description and data processing tools are as follows:
  • Application: Global context accessible across all MAS instances. Data tools are: oxy_request.get/set_global_data()
  • Session Group: Shared memory among a cluster of related conversations. Data tools are: oxy_request.get/set_group_data()
  • Request: Transient data restricted to a single inference trajectory. Data tools are: oxy_request.get/set_shared_data()
  • Node: Local arguments specific to an individual Oxy component. Data tools are: oxy_request.get/set_arguments()
Through hierarchical data isolation and sharing, the efficiency of data management and the ease of development in multi-agent collaborative scenarios are significantly improved.
3.2 Permission-Driven Planning and Standardized Execution Lifecycle
In contrast to fixed workflows, OxyGent utilizes permission-driven dynamic planning to govern agent collaboration. The execution trajectory is not predefined but emerges from the authorization relationships between Oxy nodes. To support this dynamism while maintaining observability, we implement a comprehensive Oxy execution lifecycle as shown in Figure 3. Each node traverses a standardized sequence of management steps, and the example operations for each stage are as follows:
  • Pre-execution Phase: _pre_process is used for data formatting and _pre_save_data persists inputs before logic invocation.
  • Core Execution: The _execute hook triggers the actual LLM reasoning or tool usage.
  • Post-execution Phase: _post_process for data post-processing and _format_output for downstream compatibility.
By utilizing Aspect-Oriented Programming (AOP), OxyGent injects observability and security operators into these lifecycle joinpoints. This enables the system to automatically synthesize real-time execution visualizations of the actual call graphs during runtime (Figure 4). Furthermore, as shown in Figure 5, OxyGent incorporates production-grade time tracking to expose real-time agent-level resource consumption. By decomposing execution into LLM inference, tool and API calls, and inter-agent interactions, it enables efficient bottleneck identification and fine-grained architectural optimization.
3.3 Evolutionary AI Asset Management via OxyBank
The long-term value of an MAS lies in its ability to self-improve. We introduce OxyBank, an all-in-one AI asset management platform designed to serve as the framework’s memory bank (Figure 6). OxyBank bridges the gap between raw execution data and the continuous self-improvement of the collective agent ecosystem.
The framework supports autonomous evolution by establishing a closed-loop cycle for annotation, auditing, and data backflow. Online execution traces are automatically captured and precipitated as memory assets within OxyBank. These assets are then processed by specialized annotation agents or human experts using customized templates (e.g., QA or business-domain tagging). Once audited, the high-quality samples are used to update knowledge bases or fine-tune models through joint evolution. This pipeline ensures that the MAS continuously refines its decision-making strategies based on validated feedback from the production environment.
Figure 8: A hierarchical MAS of 2,000+ agents for e-commerce classification, employing a top-down decision chain with dynamic logic and full-chain tracing. “>>” denotes a call invocation between Oxy nodes.
4 Case Studies
4.1 Evaluations on the GAIA Benchmark
Benchmark. GAIA Mialon et al. (2023) focuses on real-world tasks that require fundamental human-level abilities. It is intuitive for humans, who achieve an average accuracy of 92%, while frontier models like GPT-4, equipped with plugins, struggled at 15%. The benchmark comprises 466 questions across 3 levels of increasing complexity.
Experimental Results. To evaluate the orchestration effectiveness of OxyGent, we tested the framework against the GAIA leaderboard. On July 22, 2025, OxyGent achieved a score of 59.14%, ranking as the second-highest open-source method globally at the time, closely trailing the open-source benchmark OWL++ Hu et al. (2025) (60.8%), as shown in Figure 7. Although the emergence of newer state-of-the-art models and advanced reasoning toolsets has since moved the performance ceiling, these results highlight OxyGent’s inherent structural advantages in managing long-chain interactions and multi-modal dependencies efficiently. By leveraging its unified atomic abstraction, the framework enables high performance even with earlier-generation base models. The method is detailed in Appendix A.2. The full implementation and specific configurations used for this evaluation are available for community review at: https://github.com/jd-opensource/OxyGent/tree/gaia59.
4.2 Case Analysis in Business Scenarios
Beyond academic benchmarks, OxyGent has demonstrated significant industrial impact through wide-scale deployment in enterprise production and open-source communities. To illustrate the framework’s capacity for extreme scale and complexity, we implemented a hierarchical multi-agent classification system for e-commerce customer service, as shown in Figure 8. This system comprises over 2,000 agents tasked with categorizing incoming requests into a taxonomy of more than 2,400 distinct labels. Each category contains as few as 10 samples, making the task a massive few-shot classification problem. The architecture employs a top-down decision chain with dynamic organizational logic and full-chain tracing. Crucially, the system features a topological self-evolution mechanism that autonomously discovers an average of 5.4 new categories per week while retiring obsolete ones. In live production, this MAS-driven approach improved classification accuracy from 61.3% for single-agent-based systems to 85.6%, proving the framework’s robustness in high-concurrency, evolving business environments. Due to space constraints, additional industrial scenarios including automated SOPs, RAG-based assistants, and community-contributed tools are detailed in Appendix A.3.
5 Conclusion
In this paper, we present OxyGent, an open-source framework for building scalable and self-improving multi-agent systems. By adopting a unified atomic abstraction, OxyGent enables a Lego-like construction process that replaces rigid workflows with permission-driven dynamic planning, offering high observability through real-time execution visualizations. Integrated with OxyBank for AI asset management, the framework establishes a robust pipeline for continuous agent evolution, as evidenced by our extensive evaluations on the GAIA benchmark and industrial applications.
Limitations.
Although OxyBank supports self-evolution, large-scale training currently depends on manual resource configuration. We are actively developing intelligent resource scheduling to achieve a fully automated lifecycle, including agent construction, deployment, and refinement. This will minimize human intervention and further accelerate the evolution of collective intelligence in complex environments.
Appendix A Appendix
A.1 More Features of OxyGent
Beyond the core orchestration and evolution pipelines, OxyGent provides a suite of tools designed for deep inspection and real-time intervention in AI decision-making.
Interactive Decision Exploration. OxyGent allows researchers to pause execution at any stage to analyze an agent’s internal state, including LLM prompts, memory snapshots, and pending tool invocations. Developers can modify variables and re-sample inference paths on-the-fly without restarting the entire system. This enables parallel experimentation with different models or prompts, in which each variant is automatically recorded for seamless comparison and rollbacks, as shown in Figure 9.
Cognitive Transparency. To address the black-box nature of multi-agent systems, OxyGent provides end-to-end visibility from high-level strategies to atomic tool operations. As shown in Figure 10, the system automatically constructs traceable decision graphs that capture the rationale behind each action. This Git-like versioning of agent reasoning enables transparent collaboration and allows intermediate nodes to be regenerated under modified configurations to verify the system’s adaptability.
Figure 9: During runtime, the calling nodes on the left will be highlighted. When the inference is complete, any intermediate step can be retraced using the progress button in the upper right corner.
Figure 10: OxyGent can automatically generate traceable decision graphs in real time. It also supports configuration modification and regeneration of any intermediate node, making the MAS more transparent and operable.
A.2 Details of OxyGent-GAIA
The OxyGent-GAIA is built based on DeepSeek-R1 Guo et al. (2025), GPT-4o Hurst et al. (2024), and Claude-3.5-Sonnet Anthropic (2024). The agentic hierarchy and tool distribution are detailed below:
  • Master Agent (DeepSeek-R1): The root orchestrator that initiates and governs the entire workflow, managing the following three primary functional units:
    • Task Agent (GPT-4o): Triggered first by the Master to perform high-level task decomposition and generate the overall strategic plan for the request.
    • Coordinator Agent (GPT-4o): Manages the specific coordination and execution phase by delegating sub-tasks to its specialized sub-agents:
        Web Agent (GPT-4o): Equipped with the SearchToolkit (Google, Wikipedia, revisions, and archives), DocumentProcessingToolkit, AsyncBrowserToolkit, and VideoAnalysisToolkit for real-time information retrieval.
        Document Processing Agent (GPT-4o): Specialized in multimodal content extraction using toolkits for documents, images, audio, video, and general code execution.
        Reasoning Coding Agent (GPT-4o): Focused on complex logic and data manipulation. It utilizes the CodeExecutionToolkit, powered specifically by Claude-3.5-Sonnet, along with the ExcelToolkit and DocumentProcessingToolkit.
    • Answerer Agent (GPT-4o): Responsible for the final response synthesis. Due to GAIA’s strict formatting requirements, this agent ensures precision and has the authority to reject and re-delegate the task if the reasoning evidence is deemed insufficient.
A.3 Extended Case Studies
A.3.1 Commercial Application Patterns
The OxyGent framework has established several standardized patterns for industrial MAS deployment:
  • Standard Operating Procedures (SOP): Agents are assigned to specific workflow segments (e.g., approval, data validation). Top-level agents monitor progress while lower-level agents execute tasks, improving efficiency in standardized business processes like contract auditing.
  • Advanced RAG Pipelines: A multi-layered collaboration where specialized agents handle retrieval, answer generation, and quality review separately, significantly reducing hallucinations in enterprise knowledge bases.
  • Automated Data Analytics: OxyGent orchestrates agents to perform end-to-end data collection, cleaning, modeling, and visualization, streamlining the generation of automated business reports and anomaly detection.
  • Cross-System Tool Orchestration: In complex DevOps and office automation scenarios, agents dynamically select and invoke external APIs, maintaining system state across multiple platforms.
A.3.2 Community and Developer Feedback
Open-source community feedback highlights the extensibility of OxyGent across diverse niche domains:
  • Natural Language to SQL (NL2SQL): Developers have utilized OxyGent to build multi-agent clusters that convert natural language queries into complex SQL statements with multi-dimensional consistency verification.
  • Low-Code Automation: Integration with tools like JoyCode has enabled users to perform web searches and file operations via natural language.
  • Database Optimization: Autonomous agents built on OxyGent are currently used to diagnose “slow SQL” queries and generate governance suggestions to improve system performance.
  • Dynamic Flowchart Generation: Using local API calls, agents can automatically generate and render core implementation logic into visual flowcharts for better developer communication.